# How Books Tell a History of Statistics in Portugal: Works of Foreigners, $\boldsymbol{Estrangeirados}$, and Others

### v1

@@ 1 Introduction
The third posthumous edition of The Doctrine of Chances by Abraham de Moivre (1756) contains applications of Probability Theory to actuarial calculations regarding the computation of annuities. Daniel Augusto da Silva addressed the calculation of annuities in 1866, thus becoming the patron of Portuguese actuaries. It is an interesting investigation but of little statistical interest, the reference to statistics is limited to the excerpt we reproduce.
@@
    The values of the amortization taxes given by Statistics and formula (C) for the two Montepio Mutualist Associations that we analyse fall apart, in opposite directions, from the values they would assume in those associations, in case they would be in normal circunstances.
@@
Daniel Augusto da Silva (1870) also published a comparative study of population movements in Portugal. Detailed information can be found in Martins (2012) [18], with complements in Martins (2020) [19].
  • The Geographia e Estatística Geral de Portugal e Colónias [Geography and General Statistics of Portugal and the Colonies] by Gerardo A. Pery, from 1875, is an interesting synthesis that, without methodological innovations, plays a pioneering role in the association of Geography with Statistics. Population is also a major theme in this work, as far as statistics is concerned.
@@
  • From 1871 to 1899 Manuel Nunes Giraldes was the lecturer in Economics and Statistics at the Faculty of Law in Coimbra; in addition to José Frederico Laranjo’s collaboration, M. Emydgio Garcia gave some lectures on the course. Emydgio Garcia published the corresponding notes in 1894, where Statistics was confined to the title of the pamphlet and a set of general comments, which seems to indicate some backlashes in the teaching of Political Economy at the Faculty of Law.
@@
Other pioneering works in Mathematical Statistics, notably the posthumous publication where Thomas Bayes (1763) [22] used Probability as the language of Statistics for the first time, the great memoir of Laplace’s (1774) [23] on inference, and Gauss’ (1809) [24] research in which the method of least squares is developed, only came to have an impact on statistical production in Portugal at the end of the 19th century.
  • The first doctoral thesis at the Faculty of Mathematics of the University of Coimbra in the field of mathematical statistics, Introdução à Teoria dos Erros das Observações [Introduction to the Theory of Observational Errors], was defended in 1898 by Sidónio Paes. It proposes a rigorous demonstration of the approximation of sums by Gauss’s law, and of the use of “lesser squares”. It explicitly mentions the work of Bayes, Laplace, Lagrange and Gauss, but without precise referencing.
@@
At the end of the century, Pedro Afonso André de Figueiredo, honoured with the title of Viscount of Wildik, published Notice Statistique sur le Portugal for the Universal Exhibition of 1900.
@@
Chronological List of XIX-th Century Publications in Statistics(7)(7)(7)  In red are the institutions where the rarest works can be found: FLUL – Faculdade de Letras da Universidade de Lisboa; FDUL – Faculdade de Direito da Universidade de Lisboa; FMUL – Faculdade de Medicina da Universidade de Lisboa; IICT – Instituto de Investigação Científica Tropical; IST – Instituto Superior Técnico; ISA – Instituto Superior de Agronomia; FPCEUL – Faculdade de Psicologia e Ciências da Educação da Universidade de Lisboa; FCUL – Faculdade de Ciências da Universidade de Lisboa; ISCSP – Instituto Superior de Ciências Sociais e Políticas; ISEG – Instituto Superior de Economia e Gestão.
Solano Constâncio, F. (1818). Discurso Preliminar [Preliminary Speech], in Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo I, 1–37. (Reproduced in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, p. 81–99, available at https://www.bportugal.pt/sites/default/ files/ocpep-11.pdf).
Xavier, C. (1820). Considerações sobre a Statistica [Considerations on Statistics], Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo X, p. 134–172. http://books.google.com/books?id=6svQuAvOksoC&printsec=frontcover&dq=editions:04b0YnI0-8bkocSqJH#PRA2-PA134,M1
@@
Balbi, A. (1822). Essai statistique sur le Royaume de Portugal et d’Algarve, comparé aux autres états de l’Europe, et suivi d’un coup’oeil sur l’État actuel des sciences, des lettres et des beaux-arts parmi les portugais des deux hémishéres, Reych et Gravier, libraires, Paris.
https://books.google.fr/books?id=_Jo2AAAAMAAJ&printsec=frontcover#v=onepage&q&f=false (There is a modern edition, Imprensa Nacional – Casa da Moeda, Lisboa, 2004.)
Balbi, A. (1822a). Variétés politiques et statistiques de la monarchie portugaise, Rey et Gravier, Imp. de Casson, Paris. (reedição: Nabu Press, 2010)
(https://books.google.pt/books/about/Vari%C3%A9t%C3%A9s_politico_statistiques_sur_la.html?id =KmrIByr7oocC&redir_esc=y)
Solano Constâncio, F. (1821). An inquiry concerning population, etc…, ou investigação acerca da população e da faculdade de multiplicação da espécie humana; obra destinada a refutar a doutrina do ensaio de M. Malthus sobre esse assunto por W. Godwin, Londres, 1820 [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr.  Malthus’s essay on this subject by W. Godwin, London, 1820], Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo XII, 63–104. (Reproduced in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, pp. 171–192, available at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.)
Solano Constâncio, F. (1822). Variétés politico-statistiques sur la monarchie portugaise, por A. Balbi, Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo XVII, 84–103. (Reproduced in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, pp. 212–222, available at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.)
@@
Villela da Silva, L. D. (1828). Observações críticas sobre alguns artigos do Ensaio estatístico do reino de Portugal e Algarves, publicado em Paris por Adriano Balbi [Critical observations on some considerations in Variétés politiques et statistiques de la monarchie portugaise, published in Paris by Adran Balbi], Imprensa Régia, Lisboa. FLUL
Solano Constâncio, F. (1836). Novo Diccionario Critico e Etymologico da Lingua Portugueza [New Critical and Etymological Dictionary of the Portuguese Language], Angelo Francisco Carneiro, Paris. (Available at https://books.google.pt/books?id=HM9TAAAAcAAJ&printsec=frontcover&hl=pt-BR&source=gbs_ge_summary_r&cad=0#v=onepage&q&f=false)
Solano Constâncio, F. (1838). Considerações sobre a estatística [Considerations on Statistics], Armazém de Conhecimentos Úteis, nas Artes e Ofícios, ou Colecção de Tratados, Receitas e Invenções de Utilidade Geral [Repository of Useful Knowledge, in Arts and Crafts, or Collection of Treatises, Recipes, and Inventions of General Utility], J.-P. Aillaud, Paris, 1 (1938), 51–64. (Reprinted in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, p. 271–277, available at https://www.bportugal.pt/sites/ default/files/ocpep-11.pdf.)
Forjaz de Sampaio, A. P. (1841). Primeiros Elementos da Sciencia da Statistica [First Elements of the Science of Statistics], Imprensa de Trovão & Companhia, Coimbra. New edition (together with Elementos de Economia Política [Elements of Political Economy], Coimbra, 1839), entitled Elementos de Economia Política e Estadística [Elements of Political Economy and Statistics], Coimbra, 1845; new edition in 1852; the 1856 edition was entitled Novos Elementos de Economia Política e Estadística [New Elements of Political Economy and Statistics]; in 1858 a new edition in three volumes has been released, entitled Elementos de Economia Política [Elements of Political Economy]. Further infomation in https://www.arqnet.pt/dicionario/sampaioapf.html. Recent edition: Forjaz de Sampaio, A. (1995), Estudos e Elementos de Economia Política [Studies and Elements of Political Economy], Banco de Portugal, Lisboa with and Introduction by Alcino Pedrosa (p. XIII–XXVIII). In Tomo I, available at https://www.bportugal.pt/sites/default/files/ocpep-10_t1.pdf, reprint of Primeiros Elementos da Ciência da Estatística [First Elements of the Science of Statistics], p. 79–112. In Tomo II, available at https://www.bportugal.pt/sites/default/files/ocpep-10_t2.pdf, reprint of the 1974 edition of Elementos de Economia Política e Estadística [Elements of Political Economy and Statistics], with Primeiras Noções da Teoria da Estadística [First Elements of the Science of Statistics] reprinted on pp. 53–60, with notes 291–332 in pp. 373–420.
Santos Cruz, F.  I. (1841). Da prostituição na cidade de Lisboa [On the protitution at Lisbon], Typ. Lisbonense, Lisboa (reprinted in 1984, D. Quixote, Lisboa).
Jonnès, A. M. (1851). Compendio Estadistico ou Resumo dos Elementos de Estadistica [A Treatise on Statistics, or A Resume of the Elements of Statistics], Imprensa Nacional, Lisboa. (2ªed., de 1856, available at https://books.google.tt/books?id=OOoOAAAAQAAJ&printsec=frontcover#v =onepage&q&f=false). ISA
@@
Marreca, A. O. (1854). Parecer e memória sobre a proposta, que apresentou o Sr. Alexandre Herculano: para que a secção de Ciências Económicas e Administrativas redigisse um Projecto de Estatística [A Review and Memoir on Mr. Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Statistics], Academia Real das Ciências, Lisboa. FLUL
Ávila, A. J. (1854). Relatório sobre os Trabalhos do Congresso de Estatística reunido em Bruxelas em 1853 [Report on the Works of the Congress of Statistics held in Brussels in 1853], Imprensa Nacional, Lisboa.
Ferreira, J. D. (1857). Ensaio sobre os Primeiros Elementos da Theoria da Estadistica [Essay on the First Elements of Statistical Theory], Imprensa da Universidade, Coimbra. FDUL
@@
Ávila, A. J. (1864). Relatório sobre os trabalhos do Congresso Internacional de Estatísticas reunido em Berlin em 1863 [Report on the Works of the Congress of Statistics held in Berlin in 1863], Imprensa Nacional, Lisboa. FDUL
@@
Azevedo, F. P. (1864). História da prostituição e policia sanitária no Porto: seguida de um ensaio estatistico dos dous ultimos annos, tabelas comparativas, etc. [History of prostitution and sanitary police at Porto, followed by a statistical essay on the two elapsed years, comparative tables, etc.], F. Gomes da Fonseca, Porto. FMUL
Silva, D. A. (1866). Amortização annual media das pensões nos principaes montepios de sobrevivencia portuguezes [Average yearly amortizations at the foremost survival Montepio Mutualist Associations in Portugal], Jornal de Sciencias Matematicas, Physicas e Naturaes, Tomo I, 175–187. https://www.marinha.pt/pt/a-marinha/historia/personalidades-ilustres/Documents/Daniel%20Aug usto%20da%20Silva/05A_Amortizacao_Annual_Media_Pensoes.pdf
Silva, D. A. (1870). Contribuições para o estudo comparativo do movimento da população em Portugal [Contributions for the comparative study of the population movements in Portugal], Jornal de Sciencias Mathematicas, Physicas e Naturaes, tomo II, VIII, 255–306. FDUL
Pery, G. A. (1875). Geographia e Estatística Geral de Portugal e Colónias [Geography and General Statistics of Portugal and the Colonies], Imprensa Nacional, Lisboa. https://ia800708.us.archive.org/ 19/items/geographiaeesta02perygoog/geographiaeesta02perygoog.pdf IICT
@@
Laranjo, J. F. (1887). Programma para a cadeira de economia política e estatística da Faculdade de Direito da Universidade de Coimbra [Program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra], Imprensa da Universidade, Coimbra. FLUL
@@
Garcia, M. E (1894). Apontamentos de Algumas Prelecções no Curso de Economia Política e Estatística [Notes of Some Magistral Lectures in the Course of Political Economy and Statistics], Typ. de Luiz Cardoso, Coimbra.
Paes, S. (1898). Introdução à Teoria dos Erros das Observações [Introduction to the Theory of Observational Errors]. Imprensa da Universidade, Coimbra. https://am.uc.pt/item/44025
Wildik, Vicomte de (1900). Notice Statistique sur le Portugal, Aillaud, Paris. FLUL
@@
Karl Pearson, William Sealy Gosset (Student), Ronald Fisher, Jerzy Neyman, and Egon Pearson were major figures in the early developements of Mathematical Statistics, namely in what regards inference about parameters assuming normality, discriminant analysis, and experimental design. These concepts were gradually imported, especially through lecture notes (some published as books) by Varennes e Mendonça at ISA (Instituto Superior de Agronomia), and by Leite Pinto and by Bento Murteira at ISCEF (Instituto Superior de Ciências Económicas e Financeiras). On the other hand, the new methods of Statistics were addressed in a series of publications on “methodological issues” by W. L. Stevens (a disciple of Fisher who worked in the Anthropology Institute of Coimbra between 1942 and 1945), and by his patron E. Tamagnini (1937), and by agronomists influenced by Fisher’s school.
Outside the university context, Guimarães (1931) used statistical tests to investigate public education issues, and Amaro Guerreiro (1947)(8)(8)(8)  An interesting analysis of this book can be found at https://www.deepdyve.com/lp/wiley/manual-de-estat-stica-vwDem3M8IE?utm_source=freeShare&utm_medium=link&utm_campaign=freeShare., a senior civil servant at INE (Instituto Nacional de Estatística) included in his book a description of Pearson distributions (inserts between pp. 246 and 247) and in Chapter 10 (Curve Fitting) in the final section addresses the question of “Fitting to Pearson Type I Curve”; it also contains a chapter (8) on Index Numbers and another (9) on Time Series Analysis.
@@
Sidónio Paes’s (1898) thesis, presented at the Faculty of Mathematics of Coimbra University, in a way marked the end of the hegemony of the Faculty of Law at the University of Coimbra regarding the university teaching of Statistics in Portugal.
@@
Since 1865, a chair of Infinitesimal Calculus including the elementary teaching of Probability Calculus had been taught at the Polytechnic School of Lisbon, and in 1911 a quarterly chair of Probability Calculus was created, which in 1932 became an annual chair, taught by Victor Hugo de Lemos and Pedro Braumann (cf. Fontes de Sousa, 2005 [8]). The lecture notes from Victor Hugo de Lemos’s course was edited in 1945 by J. Cabral Madeira.
@@
  • As far as publications in the area of Probability are concerned, the first noteworthy event was the publication of a booklet intended to the layman, in 1904, volume 223 of the Biblioteca do Povo e das Escolas [Library for the People and the Schools], Noções sobre Cálculo das Probabilidades, Theoria dos Erros e Méthodo dos Mínimos Quadrados [Notions of Probability Calculus, Theory of Errors and the Least Squares Method], by Rodolpho Guimarães.
As Rui Santos (2008a, 2024a, 2024b) [26, 27, 28] comments, Rodolpho Guimarães (9)(9)(9)  Rodolfo Guimarães (1866-1918) was an officer in the Portuguese army (Colonel) and a professor at the Army School of Lisbon (Military Academy). He also served at the Astronomical Observatory of Tapada in Lisbon. Currently, Guimarães is best known for his research in the History of Mathematics, particularly for publishing a catalog of mathematical works by Portuguese authors during the 19th century and for promoting the work of Pedro Nunes (1502–1578), as well as some contributions in the field of Geometry. For detailed information on Rodolfo Guimarães, see Saraiva, L. (1997) [25]. , in the scant 64 pages of this small volume, manages to present an interesting introduction to the topic.
The second notable publication is by José Freire de Sousa Pinto (1855-1911) — the Noções de Cálculo das Probabilidades para o Estabelecimento das Bases da Estatística [Notions of Probability Calculus for Establishing the Bases of Statistics], analysed in Santos (2008b, 2016) [29, 30]. José de Sousa Pinto, a professor at the University of Coimbra, discusses the use of Probability Calculus in modeling random phenomena. In this work, the author begins by explaining the process of constructing knowledge in any science, concluding that the advancement of knowledge results from the combination of observing the phenomenon and theoretical reasoning. Thus, any theory arises from the observation of a phenomenon, initially being an approximation. Then, the phenomenon is observed once more. This new observation serves to test the initial theory, which is confirmed or refined with new knowledge from the new observation, or even altered if the initial theory is found to be erroneous. It is in this successive chain of dual effort, observation and reasoning, that scientific knowledge advances towards a complete understanding of the phenomenon, where the formalized theory will always occur whenever the phenomenon happens.
It should be noted that the analysis of prerequisites for teaching Statistics was a subject of reflections by Leite Pinto (1945), in Noções de Matemática necessárias ao estudo da Estatística [Notions of Mathematics Necessary for the Study of Statistics], and by Amaro Guerreiro, in Em Torno do Ensino da Estatística [About the Teaching of Statistics], 1952-53.
@@
  • In 1914, Diogo Pacheco de Amorim (1888-1976) defended his doctoral dissertation at the University of Coimbra on Elementos de Cálculo das Probabilidades [Elements of Probability Calculus], the aim of which was to establish rigorous foundations for the theory of Probability. In fact, this objective failed, and another doctoral thesis on Probability, by Manuel Reis (1929), is very critical of Pacheco de Amorim’s thesis.
@@
Later, in a unique work by Luís de Albuquerque (1959) on the foundations of the Calculus of Probabilities, those dissertations are not even mentioned. Pacheco de Amorim’s thesis has been exhaustively studied by Rui Santos (2008c) [31].
See, as a simple introduction, Pestana and Velosa (2001) [32] and Pestana, D. (1994) [33]. There is an annotated English translation, Mendonça et al. (2007) [34].
In fact, it was only in 1933 that Kolmogorov (1933) [35] made an axiomatic construction of Probability, inspired by Maurice Fréchet’s ideas on the use of Measure Theory.
@@
  • The Theory of Measurement was tackled by José Joaquim Dionísio (1957) (who, in what regards Probability, published a brief note on Shannon’s entropy in 1959 in Gazeta de Matemática), and became the topic of choice for Pedro Braumann (1969, 1978).
@@
Pedro Braumann, who was a fellow at Stanford in 1955-56, was also the author of an excellent exposition on the central theme of classical probability theory, the arithmetic of probability laws, [Introduction to the Study of Limits of Sums of Independent Casual Variables] (1958).

### v2

@@ 1 Introduction
The third posthumous edition (1756) of The Doctrine of Chances by Abraham de Moivre [40] contains applications of Probability Theory to actuarial calculations regarding the computation of annuities. Daniel Augusto da Silva addressed the calculation of annuities in 1866, thus becoming the patron of Portuguese actuaries. It is an interesting investigation but of little statistical interest, the reference to statistics is limited to the excerpt we reproduce.
@@
    The values of the amortization taxes given by Statistics and formula (C) for the two Montepio Mutualist Associations that we analyse fall apart, in opposite directions, from the values they would assume in those associations, in case they would be in normal circumstances.
@@
Daniel Augusto da Silva also published a comparative study of population movements in Portugal [161]. Detailed information can be found in Martins [105], with further complements in Martins [106].
  • The Geographia e Estatística Geral de Portugal e Colónias [136] [Geography and General Statistics of Portugal and the Colonies] by Gerardo A. Pery, from 1875, is an interesting synthesis that, without methodological innovations, plays a pioneering role in the association of Geography with Statistics. Population is also a major theme in this work, as far as statistics is concerned.
@@
  • From 1871 to 1899 Manuel Nunes Giraldes was the lecturer in Economics and Statistics at the Faculty of Law in Coimbra; in addition to José Frederico Laranjo’s collaboration, M. Emydgio Garcia gave some lectures on the course. Emydgio Garcia published the corresponding notes in 1894 [55], where Statistics was confined to the title of the pamphlet and a set of general comments, which seems to indicate some backlashes in the teaching of Political Economy at the Faculty of Law.
@@
Other pioneering works in Mathematical Statistics, notably the posthumous publication where Thomas Bayes (1763) [16] used Probability as the language of Statistics for the first time, the great memoir of Laplace’s (1774) [87] on inference, and Gauss’ research (1809) [56] in which the method of least squares is developed, only came to have an impact on statistical production in Portugal at the end of the 19th century.
  • The first doctoral thesis at the Faculty of Mathematics of the University of Coimbra in the field of Mathematical Statistics, Introdução à Teoria dos Erros das Observações [129] [Introduction to the Theory of Observational Errors], was defended in 1898 by Sidónio Paes. It proposes a rigorous demonstration of the approximation of sums by Gauss’s law, and of the use of “lesser squares”. It explicitly mentions the works of Bayes, Laplace, Lagrange and Gauss, but without precise referencing.
@@
At the end of the century, Pedro Afonso André de Figueiredo, honoured with the title of Viscount of Wildik, published Notice Statistique sur le Portugal [203] for the Universal Exhibition of 1900.
Chronology
@@
(English translation of Portuguese titles inside square brackets. Indication of libraries where rare books were found inside angle brackets: $<FLUL>$ – Faculdade de Letras da Universidade de Lisboa; $<FDUL>$ – Faculdade de Direito da Universidade de Lisboa; $<FMUL>$ – Faculdade de Medicina da Universidade de Lisboa; $<IICT>$ – Instituto de Investigação Científica Tropical; $<IST>$ – Instituto Superior Técnico; $<ISA>$ – Instituto Superior de Agronomia; $<FPCEUL>$ – Faculdade de Psicologia e Ciências da Educação da Universidade de Lisboa; $<FCUL>$ – Faculdade de Ciências da Universidade de Lisboa; $<ISCSP>$ – Instituto Superior de Ciências Sociais e Políticas; $<ISEG>$ – Instituto Superior de Economia e Gestão.)
1818 — F. Solano Constâncio’s [Preliminary Speech] [166].
1820 — C. Xavier’s [Considerations on Statistics] [207].
1821 — F. Solano Constâncio’s [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr.  Malthus’s essay on this subject by W. Godwin, London, 1820], [167].
@@
1822 — A. Balbi’s Essai statistique sur le Royaume de Portugal et d’Algarve, comparé aux autres états de l’Europe, et suivi d’un coup’oeil sur l’État actuel des sciences, des lettres et des beaux-arts parmi les portugais des deux hémishéres [8].
— A. Balbi’s Variétés politiques et statistiques de la monarchie portugaise [9].
— F. Solano Constâncio’s [Variétés politico-statistiques sur la monarchie portugaise, by A. Balbi] [168].
@@
1828 — L. D. Villela da Silva’s [Critical observations on some considerations in Variétés politiques et statistiques de la monarchie portugaise, published in Paris by Adran Balbi] [204]. $<FLUL>$
1836 — F. Solano Constâncio’s [New Critical and Etymological Dictionary of the Portuguese Language] [169].
— Passos Manuel’s reform of the University, namely creating of a Chair of Political Economy at the Faculty of Law of the University of Coimbra.
1838 — F. Solano Constâncio’s [Considerations on Statistics] [170].
1841 — A. P. Forjaz de Sampaio’s [First Elements of the Science of Statistics] [52].
— F. I. Santos Cruz’s [On the Prostitution at Lisbon] [156].
@@
1854 — A. Oliveira Marreca’s [A Review and Memoir on Mr. Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Stadistics] [103]. $<FLUL>$
— A. J. d’Ávila’s [Report on the Works of the Congress of Statistics held in Brussels in 1853] [5].
1856 — A. Moreau de Jonnès’ Compendio Estadistico ou Resumo dos Elementos de Estadistica [112], translation of Éléments de Statistique: comprenant les principes généraux de cette science, et un aperçu historique de ses progrès [111]. $<ISA>$
1857 — J. Dias Ferreira’s [Essay on the First Elements of Stadistics Theory] [47]. $<FDUL>$
1864 — creation of the Direção-Geral de Estatística [General Direction of Statistics].
@@
— A. J. d’Ávila’s [Report on the Works of the Congress of Statistics held in Berlin in 1863], [6]. $<FDUL>$
@@
— F. P.  Azevedo’s [History of prostitution and sanitary police at Porto, followed by a statistical essay on the two elapsed years, comparative tables, etc.] [7]. $<FMUL>$
1865 — The teaching of elementary Probability Calculus was included in the program of Infinitesimal Calculus, at the Escola Politécnica de Lisboa [Lisbon Polytechnic School].
1866 — D. A. da Silva’s [Average yearly amortizations at the foremost survival Montepio Mutualist Associations in Portugal][160].
1870 — D. A. da Silva’s [Contributions for the comparative study of the population movements in Portugal][161]. $<FDUL>$
1875 — G. A. Pery’s [Geography and General Statistics of Portugal and the Colonies] [136]. $<IICT>$
@@
1887 — J. F. Laranjo’s [Program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra] [88]. $<FLUL>$
@@
1894 — M. E. Garcia’s [Notes of Some Magistral Lectures in the Course of Political Economy and Statistics] [55].
1898 — Sidónio Paes’ [Introduction to the Theory of Observational Errorss] [129].
1900 — Vicomte de Wildik’s [Notice Statistique sur le Portugal]. [203] $<FLUL>$
@@
Karl Pearson, William Sealy Gosset (Student), Ronald Fisher, Jerzy Neyman, and Egon Pearson were major figures in the early developments of Mathematical Statistics, namely in what regards inference about parameters assuming normality, discriminant analysis, and experimental design. These concepts were gradually imported, especially through lecture notes (some published as books) by Varennes e Mendonça at ISA (Instituto Superior de Agronomia), and by Leite Pinto and by Bento Murteira at ISCEF (Instituto Superior de Ciências Económicas e Financeiras). On the other hand, the new methods of Statistics were addressed in a series of publications on “methodological issues” by W. L. Stevens (a disciple of Fisher who worked in the Anthropology Institute of Coimbra between 1942 and 1945), and by his patron E. Tamagnini, and by agronomists influenced by Fisher’s school.
Outside the university context, Guimarães [74] used statistical tests to investigate public education issues, and Amaro Guerreiro [72]777  An interesting analysis of this book can be found at https://www.deepdyve.com/lp/wiley/manual-de-estat-stica-vwDem3M8IE?utm_source=freeShare&utm_medium=link&utm_campaign=freeShare., a senior civil servant at INE (Instituto Nacional de Estatística) included in his book a description of Pearson distributions (inserts between pp. 246 and 247) and in Chapter 10 (Curve Fitting) in the final section addresses the question of “Fitting to Pearson Type I Curve”; it also contains a chapter (8) on Index Numbers and another (9) on Time Series Analysis.
@@
Sidónio Paes’s 1898 thesis [129], presented at the Faculty of Mathematics of Coimbra University, in a way marked the end of the hegemony of the Faculty of Law at the University of Coimbra in what concerns the university teaching of Statistics in Portugal.
@@
Since 1865, a chair of Infinitesimal Calculus including the elementary teaching of Probability Calculus had been taught at the Polytechnic School of Lisbon, and in 1911 a quarterly chair of Probability Calculus was created at the Faculty of Sciences, University of Lisbon, which in 1932 became an annual chair, taught by Victor Hugo de Lemos and Pedro Braumann (cf. Fontes de Sousa, 2005 [51]). The lecture notes [92] from Victor Hugo de Lemos’s course was edited in 1945 by J. Cabral Madeira.
The five volumes of Adolpho Loureiro’s Os Portos Marítimos de Portugal e ilhas Adjacentes: Atlas [95] [The Sea Ports of Portugal and Adjacent Islands], containing many statistical tables and details, were published from 1904 to 1909.
@@
  • As far as publications in the area of Probability are concerned, the first noteworthy event was the publication of a booklet intended to the layman, in 1904, volume 223 of the Biblioteca do Povo e das Escolas [Library for the People and the Schools], Noções sobre Cálculo das Probabilidades, Theoria dos Erros e Méthodo dos Mínimos Quadrados [75] [Notions of Probability Calculus, Theory of Errors and the Least Squares Method], by Rodolpho Guimarães.
As Rui Santos [150, 154, 155] comments, Rodolpho Guimarães888  Rodolfo Guimarães (1866-1918) was a Colonel in the Portuguese army and a professor at the Army School of Lisbon (Military Academy). He also served at the Astronomical Observatory of Tapada in Lisbon. Currently, Guimarães is best known for his research in the History of Mathematics, particularly for publishing a catalog of mathematical works by Portuguese authors during the 19th century and for promoting the work of Pedro Nunes (1502–1578), as well as some contributions in the field of Geometry. For detailed information on Rodolfo Guimarães, see Saraiva [157]. , in the scant 64 pages of this small volume, manages to present an interesting introduction to the topic.
The second notable publication is by José Freire de Sousa Pinto (1855-1911) — the Noções de Cálculo das Probabilidades para o Estabelecimento das Bases da Estatística [173] [Notions of Probability Calculus for Establishing the Bases of Statistics], analysed in Santos [151, 153]. José de Sousa Pinto, a professor at the University of Coimbra, discusses the use of Probability Calculus in modeling random phenomena. In this work, the author begins by explaining the process of constructing knowledge in any experimental science, concluding that the advancement of knowledge results from the combination of observing the phenomenon and theoretical reasoning. Thus, any theory arises from the observation of a phenomenon, initially being an approximation. Then, the phenomenon is observed once more. This new observation serves to test the initial theory, which is confirmed or refined with new knowledge from the new observation, or even altered if the initial theory is found to be erroneous. It is in this successive chain of dual effort, observation and reasoning, that scientific knowledge advances towards a complete understanding of the phenomenon, where the formalized theory will always occur whenever the phenomenon happens.
It should be noted that the analysis of prerequisites for teaching Statistics was a subject of reflections by Leite Pinto [90], in Noções de Matemática necessárias ao estudo da Estatística [Notions of Mathematics Necessary for the Study of Statistics], and by Amaro Guerreiro, in Em Torno do Ensino da Estatística [73] [About the Teaching of Statistics], 1952-53.
@@
  • In 1914, Diogo Pacheco d’Amorim (1888-1976) defended his doctoral dissertation at the University of Coimbra on Elementos de Cálculo das Probabilidades [127] [Elements of Probability Calculus], the aim of which was to establish rigorous foundations for the theory of Probability. In fact, this objective failed, and another doctoral thesis on Probability [144], by Manuel Reis (1929), is very critical of Pacheco d’Amorim’s thesis.
@@
Later, in a unique work by Luís de Albuquerque [2] on the foundations of the Calculus of Probabilities, those dissertations are not even mentioned. Pacheco d’Amorim’s thesis has been exhaustively studied by Rui Santos [152], see also [107].
See, as a simple introduction to the work of Pacheco d’Amorim, Pestana and Velosa [139] and Pestana [138]. There is an annotated English translation of his doctoral thesis by Mendonça et al.  [107].
In fact, it was only in 1933 that Kolmogorov [86] made an axiomatic construction of Probability, inspired by Maurice Fréchet’s ideas on the use of Measure Theory.
@@
  • Measure Theory was tackled by José Joaquim Dionísio [43] (who, in what regards Probability, published a brief note on Shannon’s entropy [44] in Gazeta de Matemática), and became the preferred field of work for Pedro Braumann [19, 20].
@@
Pedro Braumann, who was a fellow at Stanford in 1955-56, was also the author of an excellent exposition on the central theme of classical Probability Theory, the arithmetic of probability laws, [Introduction to the Study of Limits of Sums of Independent Casual Variables] [18].